After a maximum number of 5-cm plastic cubs are packed into a rectangular box measuring 33 cm by 22 cm by 20 cm, what's the volume of the space left in the box?
Answer by stanbon(75887)

After a maximum number of 5-cm plastic cubs are packed into a rectangular box measuring 33 cm by 22 cm by 20 cm, what's the volume of the space left in the box?
---------
Volume of 1 cube = 5^3 = 125 cu in
------
Volume of the box = 33*22*20 = 14520 cu in
----
# of cubes in the box:: 14520/125 = 116
----
Ans:: 14520-116*125 = 20 cu in (amt of space left)
------------
Cheers,
Stan H.
-------------